# Computable Thermodynamic Formalism

### v1

@@ 4. Approach I
Lemma 4.2 ensures that the set $C(X)_{\phi,P_{T}}^{*}$ is well-defined for any continuous map $T$ on a compact metric space $X$ and any continuous function $\phi$. We now record a characterization of this space from [Wa92, Theorem 3 (i)].
@@
$P(T_{n},\phi_{n})-\mathopen{\langle}\mu,\phi_{n}\mathclose{\rangle}$ $>P\bigl(T_{n},c\phi+(1-c)\phi_{n}\bigr)-\mathopen{\big\langle}\mu,c\phi+(1-c)\phi_{n}\mathclose{\big\rangle}$
@@
Proof of Claim 2. We construct a sequence $\{F_{m,n}\}_{(m,n)\in\mathbb{N}^{2}}$ of uniformly computable functions such that for each $n\in\mathbb{N}$, the sequence $\{F_{m,n}\}_{m\in\mathbb{N}}$ is non-increasing and converges pointwise to $f_{n}$ as $m\to\infty$. By property (i) of Theorem 1.1, $\{P(T_{n},\psi_{n,i})\}_{(n,i)\in\mathbb{N}^{2}}$ is a sequence of uniformly upper semi-computable real numbers. Hence, by Definition 3.6, there exists an algorithm $\mathcal{A}_{p}$ such that for all $m,\,n,\,i\in\mathbb{N}$, on input m,n,im,\,n,\,i, the algorithm $\mathcal{A}_{p}$ outputs $p_{m,n,i}\in\mathcal{Q}$ such that $\{p_{m,n,i}\}_{(m,n,i)\in\mathbb{N}^{3}}$ is non-increasing in $m$ and converges to $P(T_{n},\psi_{n,i})$ as $m\rightarrow+\infty$ for all $n,i\in\mathbb{N}$. Define integral functionals $\mathcal{I}_{n,i}(\nu)\coloneqq\mathopen{\langle}\nu,\phi_{n}-\psi_{n,i}\mathclose{\rangle}$ for all $n,i\in\mathbb{N}$ and $\nu\in\mathcal{P}(X)$. Then by Proposition 3.27, it follows from the uniform computability of $\{\psi_{n,i}\}_{(n,i)\in\mathbb{N}^{2}}$ and $\{\phi_{n}\}_{n\in\mathbb{N}}$ that $\{\mathcal{I}_{n,i}\}_{(n,i)\in\mathbb{N}^{2}}$ is a sequence of uniformly computable functions. Define $F_{m,n}(\nu)\coloneqq\min\{p_{m,n,i}+\mathcal{I}_{n,i}(\nu):i\in\mathbb{N}\cap[1,m]\}$ for all $n,m\in\mathbb{N}$ and $\nu\in\mathcal{P}(X)$. By the uniform computability of $\{\mathcal{I}_{n,i}\}_{(n,i)\in\mathbb{N}^{2}}$, there exists an algorithm $\mathcal{A}_{\mathcal{I}}$ such that for all $n,\,i,\,k\in\mathbb{N}$ and $\nu\in\mathcal{P}(X)$, on input n,i,kn,\,i,\,k and an oracle for $\nu$, $\mathcal{A}_{\mathcal{I}}$ outputs $q_{i,k}$ with $\mathopen{\lvert}q_{i,k}-\mathcal{I}_{n,i}(\nu)\mathclose{\rvert}\leqslant 2^{-k}$. We can now design an algorithm $\mathcal{A}_{F}$ to compute $\{F_{m,n}\}_{(m,n)\in\mathbb{N}^{2}}$. For each $m,\,n,\,k\in\mathbb{N}$ and each $\nu\in\mathcal{P}(X)$, we apply $\mathcal{A}_{p}$ and $\mathcal{A}_{\mathcal{I}}$ to compute $\{p_{m,n,i}\}_{i=1}^{m}$ and $\{q_{i,k}\}_{i=1}^{m}$ such that $\mathopen{\lvert}q_{i,k}-\mathcal{I}_{n,i}(\nu)\mathclose{\rvert}\leqslant 2^{-k}$ for each integer $1\leqslant i\leqslant m$. Then we compute $\min\{p_{m,n,i}+q_{i,k}:i\in\mathbb{N}\cap[1,m]\}$ as the output of $\mathcal{A}_{F}$. Note that $\min\{p_{m,n,i}+q_{i,k}:i\in\mathbb{N}\cap[1,m]\}+2^{-k}\geqslant F_{m,n}(\nu)\geqslant\min\{p_{m,n,i}+q_{i,k}:i\in\mathbb{N}\cap[1,m]\}-2^{-k}$. It follows that $\mathcal{A}_{F}$ demonstrates the uniform computability of the sequence $\{F_{m,n}\}_{(m,n)\in\mathbb{N}^{2}}$ of functions.
@@
Now fix an integer $n\in\mathbb{N}$ and a measure $\nu\in\mathcal{P}(X)$. Since $\{p_{m,n,i}\}_{(m,i)\in\mathbb{N}^{2}}$ is non-increasing in $m$ for each $i\in\mathbb{N}$, the sequence $\{F_{m,n}(\nu)\}_{m\in\mathbb{N}}$ is non-increasing; so its limit exists. Note that $p_{m,n,i}\geqslant P(T_{n},\psi_{n,i})$ for all $m,\,i\in\mathbb{N}$. Then by construction, we have $F_{m,n}(\nu)\geqslant f_{n}(\nu)$ for each $m\in\mathbb{N}$. Hence, $\lim_{m\to+\infty}F_{m,n}(\nu)\geqslant f_{n}(\nu)$. On the other hand, since $\lim_{m\to+\infty}p_{m,n,i}=P(T_{n},\psi_{n,i})$ for each $i\in\mathbb{N}$, it follows from (4.6) that for each $\epsilon>0$, there exist integers $i,j$ with $1\leqslant i\leqslant j$ such that $f_{n}(\nu)+\epsilon>P(T_{n},\psi_{n,i})+\mathcal{I}_{n,i}(\nu)+\epsilon/2>p_{j,n,i}+\mathcal{I}_{n,i}(\nu)\geqslant F_{j,n}(\nu)\geqslant\lim_{m\to+\infty}F_{m,n}(\nu)$. Consequently, $f_{n}(\nu)\geqslant\lim_{m\to+\infty}F_{m,n}(\nu)$. This shows that $\lim_{m\to+\infty}F_{m,n}(\nu)=f_{n}(\nu)$.
@@
Since $\{F_{m,n}\}_{(m,n)\in\mathbb{N}^{2}}$ is a sequence of uniformly computable functions and $\{F_{m,n}(\nu)\}_{m\in\mathbb{N}}$ is non-increasing for all $\nu$, by Definition 3.15, $\{f_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly upper semi-computable functions. This establishes Claim 2.
@@
We now complete the proof by expressing $\Gamma_{n}$ as the complement of a uniformly lower semi-computable open set within the uniformly recursively compact set $\mathcal{M}(X,T_{n})$. By property (ii) of Theorem 1.1, $\{P(T_{n},\phi_{n})\}_{n\in\mathbb{N}}$ is a sequence of uniformly lower semi-computable real numbers. By Definition 3.6, there exists a sequence $\{q_{m,n}\}_{(m,n)\in\mathbb{N}^{2}}$ of uniformly computable real numbers such that, for each $n\in\mathbb{N}$, the sequence $\{q_{m,n}\}_{m\in\mathbb{N}}$ is non-decreasing in $m$ and converges to $P(T_{n},\phi_{n})$ as $m\to+\infty$. Hence, it follows from the definition of $\{D_{n}\}_{n\in\mathbb{N}}$, Claim 1, and (4.6) that
@@
Let $\Gamma_{n}$ be as defined in (4.3). Consider an arbitrary $n\in\mathbb{N}$. Then the measure-theoretic entropy map $\nu\mapsto h_{\nu}(T_{n})$ is upper semi-continuous by hypothesis. This implies that $\overline{h}_{\nu}(T_{n})=h_{\nu}(T_{n})$ for all $\nu\in\mathcal{M}(X,T_{n})$. It follows from (4.3) that the set $\Gamma_{n}$ coincides with the set of equilibrium states $\mathcal{E}(T_{n},\phi_{n})$. Hence, by property (iii), we have $\Gamma_{n}=\mathcal{E}(T_{n},\phi_{n})=\{\mu_{n}\}$.
@@
5. Approach II
@@ 5.1. Jacobian and the transfer operator
The lower bound given above is a classical result in ergodic theory known as the Rokhlin entropy formula. We refer the reader to [Sa99, Theorem 4.2] for its topological Markov shift version, [Co12, Corollary 12.1] for its finite admissible partitions version; see [He25] for the proof in our context. The uniqueness of the Jacobian immediately follows from [Ro49, Theorem 2.7], [PU10, Definition 2.9.2 & Proposition 2.9.5].
@@
    $\Psi(x)\coloneqq 0\quad\text{for each }x\in Y^{c}\quad\text{and}\quad\Psi(x)\coloneqq\Phi_{k}(T(x))\quad\text{for each }k\in\mathbb{N}\text{ and each }x\in Y_{k}.$ (5.3)
@@
(i) Fix an arbitrary $k\in\mathbb{N}$. Since $(X,\,\rho,\,T,\,Y,\,\{Y_{k}\}_{k\in\mathbb{N}},\,\mu)$ is admissible, by Definition 5.2 (ii), (iii), and (v), $T$ is Borel measurable, $Y_{k},\,T(Y_{k})\in\mathcal{B}(X)$, and $\mu\in\mathcal{M}(X,T)\cap\mathcal{P}(X;Y)$. Define $\mu_{k}(B)\coloneqq\mu(B)$ and $\widetilde{\mu}_{k}(B)\coloneqq\mu\bigl(\bigl(T|_{Y_{k}}\bigr)^{-1}(B)\bigr)=\mu\bigl(T^{-1}(B)\cap Y_{k}\bigr)$ for each Borel subset $B\subseteq T(Y_{k})$. Then $\mu_{k}$ and $\widetilde{\mu}_{k}$ are both $\sigma$-finite positive Borel measures on $T(Y_{k})$. Since $\mu\in\mathcal{M}(X,T)\cap\mathcal{P}(X;Y)$, we obtain that $\widetilde{\mu}_{k}(B)=\mu\bigl(T^{-1}(B)\cap Y_{k}\bigr)\leqslant\mu\bigl(T^{-1}(B)\bigr)=\mu(B)=\mu_{k}(B)$ for each Borel subset $B\subseteq T(Y_{k})$. This implies that $\widetilde{\mu}_{k}$ is absolutely continuous with respect to $\mu_{k}$. By the Radon–Nikodym theorem, there exists a nonnegative $\mu_{k}$-integrable derivative $\frac{\mathrm{d}\widetilde{\mu}_{k}}{\mathrm{d}\mu_{k}}$. Define $\Phi_{k}:X\rightarrow\mathbb{R}$ by
@@
Finally, we verify that $J_{\mu}(x)\cdot\Psi(x)=1$ for $\mu$-a.e. $x\in X$. By [Ke95, Corollary 15.2], it follows from Definition 5.2 (iii) that $T$ is a Borel isomorphism of $Y_{k}$ with $T(Y_{k})$ for each $k\in\mathbb{N}$. Since $J_{\mu}$ is a Jacobian on $\widetilde{Y}$ for $T$ with respect to $\mu$, by the definition of $\bigl\{\tilde{\nu}_{k}\bigr\}_{k\in\mathbb{N}}$, we have
@@
$\mu(A)=\mu\bigl(T^{-1}(T(A))\cap Y_{k}\bigr)=\int_{T(A)}\!\Phi_{k}\,\mathrm{d}\mu=\int_{A}\!(\Phi_{k}\circ T)\,\mathrm{d}(\mu\circ T)=\int_{A}\!\Psi\,\mathrm{d}\tilde{\nu}_{k}$
@@
$\int_{A}\!\mathcal{L}_{\mu}(\mathbbm{1})\,\mathrm{d}\mu$ $=\int_{A}\!\biggl(\sum_{y\in T^{-1}(x)\cap Y}\Psi(y)\biggr)\,\mathrm{d}\mu(x)=\int_{A}\!\biggl(\sum_{k\in\mathbb{N}:T^{-1}(x)\cap Y_{k}\neq\emptyset}\Phi_{k}\biggr)\,\mathrm{d}\mu=\int_{A}\!\biggl(\sum_{k\in\mathbb{N}}\Phi_{k}\biggr)\,\mathrm{d}\mu$
@@
Since $J$ is a positive Borel function and $J_{\mu}$ is a nonnegative Borel function, we have $J_{\mu}/J\in L^{+}(X)$. By (5.4), $\widetilde{Y}=Y\smallsetminus\Psi^{-1}(0)$, (5.14), and (5.12), we obtain that for $\mu$-a.e. $x\in X$,
$\mathcal{L}_{\mu}\bigl(J_{\mu}/J\bigr)(x)$ $=\sum_{y\in T^{-1}(x)\cap Y}\frac{J_{\mu}(y)\Psi(y)}{J(y)}=\sum_{y\in T^{-1}(x)\cap\widetilde{Y}}\frac{J_{\mu}(y)\Psi(y)}{J(y)}$ (5.15)
@@
Recall the definition of $\mathcal{M}(X,T;Y,J)$ from (1.2). For each admissible sextuple $(X,\,\rho,\,T,\,Y,\,\{Y_{k}\}_{k\in\mathbb{N}},\,\mu)$ and each Borel function $\phi\colon X\rightarrow\mathbb{R}$, we define
$\mathcal{E}_{0}(T,\phi)\coloneqq\bigl\{\mu\in\mathcal{E}(T,\phi)\cap\mathcal{P}(X;Y):h_{\mu}(T)=\bigl\langle\mu,\log\bigl(J_{\mu}\bigr)\bigr\rangle\bigr\},$ (5.17)
where $J_{\mu}$ is given as in Proposition 5.4 for the measure $\mu$.
@@
Then $\mu\in\mathcal{M}(X,T;Y,J)$ if and only if $\mu\in\mathcal{E}_{0}(T,\phi)$.
@@
Since $h$ is a bounded Borel function and $\mu\in\mathcal{M}(X,T)$, we have $\mathopen{\langle}\mu,h\mathclose{\rangle}=\mathopen{\langle}\mu,h\circ T\mathclose{\rangle}\in(-\infty,+\infty)$. Since $J$ satisfies property (i) of Theorem 5.6, $\mu\in\mathcal{P}(X;Y)$, and $\mathopen{\langle}\mu,\phi\mathclose{\rangle}<+\infty$, we obtain that
@@
By Proposition 5.4 (ii), there exists a Borel function $J_{\mu}\colon X\rightarrow\mathbb{R}^{+}$ which is a Jacobian for $T$ with respect to $\mu$.
We first establish the forward implication. Let $\mu\in\mathcal{M}(X,T;Y,J)$. Since $\mu\in\mathcal{P}(X;Y)$, it follows from (1.2) that $J_{\mu}(x)\geqslant J(x)$ for $\mu$-a.e. $x\in X$. By (5.18), (3.2), and Proposition 5.3, we have

### v2

@@ 4. Approach I: Global approach via convex analysis
Lemma 4.2 ensures that the set $C(X)_{\phi,P_{T}}^{*}$ is well defined for any continuous map $T$ on a compact metric space $X$ and any continuous function $\phi$. We now record a characterization of this space from [Wa92, Theorem 3 (i)].
@@
$P(T_{n},\phi_{n})-\mathopen{\langle}\mu,\phi_{n}\mathclose{\rangle}$ $>P(T_{n},c\phi+(1-c)\phi_{n})-\mathopen{\langle}\mu,c\phi+(1-c)\phi_{n}\mathclose{\rangle}$
@@
Proof of Claim 2. We construct a sequence $\{F_{m,n}\}_{(m,n)\in\mathbb{N}^{2}}$ of uniformly computable functions such that for each $n\in\mathbb{N}$, the sequence $\{F_{m,n}\}_{m\in\mathbb{N}}$ is nonincreasing and converges pointwise to $f_{n}$ as $m\to\infty$. By property (i) of Theorem 1.1, $\{P(T_{n},\psi_{n,i})\}_{(n,i)\in\mathbb{N}^{2}}$ is a sequence of uniformly upper semi-computable real numbers. Hence, by Definition 3.6, there exists an algorithm $\mathcal{A}_{p}$ such that for all $m,\,n,\,i\in\mathbb{N}$, on input m,n,im,\,n,\,i, the algorithm $\mathcal{A}_{p}$ outputs $p_{m,n,i}\in\mathcal{Q}$ such that $\{p_{m,n,i}\}_{(m,n,i)\in\mathbb{N}^{3}}$ is nonincreasing in $m$ and converges to $P(T_{n},\psi_{n,i})$ as $m\rightarrow+\infty$ for all $n,i\in\mathbb{N}$. Define integral functionals $\mathcal{I}_{n,i}(\nu)\coloneqq\mathopen{\langle}\nu,\phi_{n}-\psi_{n,i}\mathclose{\rangle}$ for all $n,i\in\mathbb{N}$ and $\nu\in\mathcal{P}(X)$. Then by Proposition 3.27, it follows from the uniform computability of $\{\psi_{n,i}\}_{(n,i)\in\mathbb{N}^{2}}$ and $\{\phi_{n}\}_{n\in\mathbb{N}}$ that $\{\mathcal{I}_{n,i}\}_{(n,i)\in\mathbb{N}^{2}}$ is a sequence of uniformly computable functions. Define $F_{m,n}(\nu)\coloneqq\min\{p_{m,n,i}+\mathcal{I}_{n,i}(\nu):i\in\mathbb{N}\cap[1,m]\}$ for all $n,m\in\mathbb{N}$ and $\nu\in\mathcal{P}(X)$. By the uniform computability of $\{\mathcal{I}_{n,i}\}_{(n,i)\in\mathbb{N}^{2}}$, there exists an algorithm $\mathcal{A}_{\mathcal{I}}$ such that for all $n,\,i,\,k\in\mathbb{N}$ and $\nu\in\mathcal{P}(X)$, on input n,i,kn,\,i,\,k and an oracle for $\nu$, $\mathcal{A}_{\mathcal{I}}$ outputs $q_{i,k}$ with $\mathopen{\lvert}q_{i,k}-\mathcal{I}_{n,i}(\nu)\mathclose{\rvert}\leqslant 2^{-k}$. We can now design an algorithm $\mathcal{A}_{F}$ to compute $\{F_{m,n}\}_{(m,n)\in\mathbb{N}^{2}}$. For each $m,\,n,\,k\in\mathbb{N}$ and each $\nu\in\mathcal{P}(X)$, we apply $\mathcal{A}_{p}$ and $\mathcal{A}_{\mathcal{I}}$ to compute $\{p_{m,n,i}\}_{i=1}^{m}$ and $\{q_{i,k}\}_{i=1}^{m}$ such that $\mathopen{\lvert}q_{i,k}-\mathcal{I}_{n,i}(\nu)\mathclose{\rvert}\leqslant 2^{-k}$ for each integer $1\leqslant i\leqslant m$. Then we compute $\min\{p_{m,n,i}+q_{i,k}:i\in\mathbb{N}\cap[1,m]\}$ as the output of $\mathcal{A}_{F}$. Note that $\min\{p_{m,n,i}+q_{i,k}:i\in\mathbb{N}\cap[1,m]\}+2^{-k}\geqslant F_{m,n}(\nu)\geqslant\min\{p_{m,n,i}+q_{i,k}:i\in\mathbb{N}\cap[1,m]\}-2^{-k}$. It follows that $\mathcal{A}_{F}$ demonstrates the uniform computability of the sequence $\{F_{m,n}\}_{(m,n)\in\mathbb{N}^{2}}$ of functions.
@@
Now fix an integer $n\in\mathbb{N}$ and a measure $\nu\in\mathcal{P}(X)$. Since $\{p_{m,n,i}\}_{(m,i)\in\mathbb{N}^{2}}$ is nonincreasing in $m$ for each $i\in\mathbb{N}$, the sequence $\{F_{m,n}(\nu)\}_{m\in\mathbb{N}}$ is nonincreasing; so its limit exists. Note that $p_{m,n,i}\geqslant P(T_{n},\psi_{n,i})$ for all $m,\,i\in\mathbb{N}$. Then by construction, we have $F_{m,n}(\nu)\geqslant f_{n}(\nu)$ for each $m\in\mathbb{N}$. Hence, $\lim_{m\to+\infty}F_{m,n}(\nu)\geqslant f_{n}(\nu)$. On the other hand, since $\lim_{m\to+\infty}p_{m,n,i}=P(T_{n},\psi_{n,i})$ for each $i\in\mathbb{N}$, it follows from (4.6) that for each $\epsilon>0$, there exist integers $i,j$ with $1\leqslant i\leqslant j$ such that $f_{n}(\nu)+\epsilon>P(T_{n},\psi_{n,i})+\mathcal{I}_{n,i}(\nu)+\epsilon/2>p_{j,n,i}+\mathcal{I}_{n,i}(\nu)\geqslant F_{j,n}(\nu)\geqslant\lim_{m\to+\infty}F_{m,n}(\nu)$. Consequently, $f_{n}(\nu)\geqslant\lim_{m\to+\infty}F_{m,n}(\nu)$. This shows that $\lim_{m\to+\infty}F_{m,n}(\nu)=f_{n}(\nu)$.
@@
Since $\{F_{m,n}\}_{(m,n)\in\mathbb{N}^{2}}$ is a sequence of uniformly computable functions and $\{F_{m,n}(\nu)\}_{m\in\mathbb{N}}$ is nonincreasing for all $\nu$, by Definition 3.15, $\{f_{n}\}_{n\in\mathbb{N}}$ is a sequence of uniformly upper semi-computable functions. This establishes Claim 2.
@@
We now complete the proof by expressing $\Gamma_{n}$ as the complement of a uniformly lower semi-computable open set within the uniformly recursively compact set $\mathcal{M}(X,T_{n})$. By property (ii) of Theorem 1.1, $\{P(T_{n},\phi_{n})\}_{n\in\mathbb{N}}$ is a sequence of uniformly lower semi-computable real numbers. By Definition 3.6, there exists a sequence $\{q_{m,n}\}_{(m,n)\in\mathbb{N}^{2}}$ of uniformly computable real numbers such that, for each $n\in\mathbb{N}$, the sequence $\{q_{m,n}\}_{m\in\mathbb{N}}$ is nondecreasing in $m$ and converges to $P(T_{n},\phi_{n})$ as $m\to+\infty$. Hence, it follows from the definition of $\{D_{n}\}_{n\in\mathbb{N}}$, Claim 1, and (4.6) that
@@
Let $\Gamma_{n}$ be as defined in (4.3). Consider an arbitrary $n\in\mathbb{N}$. Then the measure-theoretic entropy map $\nu\mapsto h_{\nu}(T_{n})$ is upper semicontinuous by hypothesis. This implies that $\overline{h}_{\nu}(T_{n})=h_{\nu}(T_{n})$ for all $\nu\in\mathcal{M}(X,T_{n})$. It follows from (4.3) that the set $\Gamma_{n}$ coincides with the set of equilibrium states $\mathcal{E}(T_{n},\phi_{n})$. Hence, by property (iii), we have $\Gamma_{n}=\mathcal{E}(T_{n},\phi_{n})=\{\mu_{n}\}$.
@@
5. Approach II: Local approach via transfer operator and Jacobian
@@ 5.1. Jacobian and the transfer operator
The lower bound given above is a classical result in ergodic theory known as the Rokhlin entropy formula. We refer the reader to [Sa99, Theorem 4.2] for a version for topological Markov shifts and to [Co12, Corollary 12.1] for a version for finite admissible partitions; see [He25, Proposition 4.3] for the proof in our context. The uniqueness of the Jacobian immediately follows from [Ro49, Theorem 2.7], [PU10, Definition 2.9.2 & Proposition 2.9.5].
@@
[…]
@@
(i) Fix an arbitrary $k\in\mathbb{N}$. Since $(X,\,\rho,\,T,\,Y,\,\{Y_{k}\}_{k\in\mathbb{N}},\,\mu)$ is admissible, by Definition 5.2 (ii), (iii), and (v), $T$ is Borel measurable, $Y_{k},\,T(Y_{k})\in\mathcal{B}(X)$, and $\mu\in\mathcal{M}(X,T)\cap\mathcal{P}(X;Y)$. Define $\mu_{k}(B)\coloneqq\mu(B)$ and $\widetilde{\mu}_{k}(B)\coloneqq\mu\bigl((T|_{Y_{k}})^{-1}(B)\bigr)=\mu\bigl(T^{-1}(B)\cap Y_{k}\bigr)$ for each Borel subset $B\subseteq T(Y_{k})$. Then $\mu_{k}$ and $\widetilde{\mu}_{k}$ are both $\sigma$-finite positive Borel measures on $T(Y_{k})$. Since $\mu\in\mathcal{M}(X,T)\cap\mathcal{P}(X;Y)$, we obtain that $\widetilde{\mu}_{k}(B)=\mu\bigl(T^{-1}(B)\cap Y_{k}\bigr)\leqslant\mu\bigl(T^{-1}(B)\bigr)=\mu(B)=\mu_{k}(B)$ for each Borel subset $B\subseteq T(Y_{k})$. This implies that $\widetilde{\mu}_{k}$ is absolutely continuous with respect to $\mu_{k}$. By the Radon–Nikodym theorem, there exists a nonnegative $\mu_{k}$-integrable derivative $\frac{\mathrm{d}\widetilde{\mu}_{k}}{\mathrm{d}\mu_{k}}$. Define $\Phi_{k}:X\rightarrow\mathbb{R}$ by
@@
Finally, we verify that $J_{\mu}(x)\cdot\Psi(x)=1$ for $\mu$-a.e. $x\in X$. By [Ke95, Corollary 15.2], it follows from Definition 5.2 (iii) that $T$ is a Borel isomorphism of $Y_{k}$ with $T(Y_{k})$ for each $k\in\mathbb{N}$. Since $J_{\mu}$ is a Jacobian on $\widetilde{Y}$ for $T$ with respect to $\mu$, by the definition of $\{\tilde{\nu}_{k}\}_{k\in\mathbb{N}}$, we have
@@
$\mu(A)=\mu\bigl(T^{-1}(T(A))\cap Y_{k}\bigr)=\int_{T(A)}\!\Phi_{k}\,\mathrm{d}\mu=\int_{A}\!\Phi_{k}\circ T\,\mathrm{d}(\mu\circ T)=\int_{A}\!\Psi\,\mathrm{d}\tilde{\nu}_{k}$
@@
$\int_{A}\!\mathcal{L}_{\mu}(\mathbbm{1})\,\mathrm{d}\mu$ $=\int_{A}\sum_{y\in T^{-1}(x)\cap Y}\Psi(y)\,\mathrm{d}\mu(x)=\int_{A}\sum_{k\in\mathbb{N}:T^{-1}(x)\cap Y_{k}\neq\emptyset}\Phi_{k}\,\mathrm{d}\mu=\int_{A}\sum_{k\in\mathbb{N}}\Phi_{k}\,\mathrm{d}\mu$
@@
Since $J$ is a positive Borel function and $J_{\mu}$ is a nonnegative Borel function, we have $J_{\mu}\big/J\in L^{+}(X)$. By (5.3), $\widetilde{Y}=Y\smallsetminus\Psi^{-1}(0)$, (5.13), and (5.11), we obtain that for $\mu$-a.e. $x\in X$,
$\mathcal{L}_{\mu}\bigl(J_{\mu}\big/J\bigr)(x)$ $=\sum_{y\in T^{-1}(x)\cap Y}\frac{J_{\mu}(y)\Psi(y)}{J(y)}=\sum_{y\in T^{-1}(x)\cap\widetilde{Y}}\frac{J_{\mu}(y)\Psi(y)}{J(y)}$ (5.14)
@@
Recall the definition of $\mathcal{M}(X,T;Y,J)$ from (1.2). For each triple $(X,\,\rho,\,T)$ satisfying properties (i) and (ii) in Definition 5.2, each Borel $Y\subseteq X$ such that there exists a unique Jacobian $J_{\mu}$ (cf. Proposition 5.3) for $T$ with respect to $\mu$ for each $\mu\in\mathcal{M}(X,T)\cap\mathcal{P}(X,Y)$, and each Borel function $\phi\colon X\rightarrow\mathbb{R}$, we define
$\mathcal{E}_{0}(T,\phi;Y)\coloneqq\bigl\{\mu\in\mathcal{E}(T,\phi)\cap\mathcal{P}(X;Y):h_{\mu}(T)=\bigl\langle\mu,\log\bigl(J_{\mu}\bigr)\bigr\rangle\bigr\},$ (5.16)
Indeed, for a Borel subset $Y\subseteq X$ satisfying that there exists $\{Y_{k}\}_{k\in\mathbb{N}}$ such that properties (iii) and (iv) in Definition 5.2 are satisfied, the existence and uniqueness of the Jacobians follow from Propositions 5.4 (ii) and 5.3. Moreover, for $Y$ with a weaker assumption, the uniqueness and existence of Jacobians still hold (cf. Lemma 5.7).
@@
Then $\mu\in\mathcal{M}(X,T;Y,J)$ if and only if $\mu\in\mathcal{E}_{0}(T,\phi;Y)$.
@@
Since $h$ is a bounded Borel function and $\mu\in\mathcal{M}(X,T)$, we have $\mathopen{\langle}\mu,h\mathclose{\rangle}=\mathopen{\langle}\mu,h\circ T\mathclose{\rangle}\in(-\infty,+\infty)$. Since $J$ satisfies property (i), $\mu\in\mathcal{P}(X;Y)$, and $\mathopen{\langle}\mu,\phi\mathclose{\rangle}<+\infty$, we obtain that
@@
By Proposition 5.4 (ii), there exists a Borel function $J_{\mu}\colon X\rightarrow\mathbb{R}^{+}$ that is a Jacobian for $T$ with respect to $\mu$.
We first establish the forward implication. Consider $\mu\in\mathcal{M}(X,T;Y,J)$. Since $\mu\in\mathcal{P}(X;Y)$, it follows from (1.2) that $J_{\mu}(x)\geqslant J(x)$ for $\mu$-a.e. $x\in X$. By (5.17), (3.2), and Proposition 5.3, we have